The fundamental frequency of a standing wave on a 1.1-m-long string is 450Hz.

What would be the wave speed of a pulse moving along this string?
Answer #1
you have to know that the fundamental wavelength is twice the length of the string, so...
wavelength = 2 * 1.1 = 2.2 meters
and
speed = wavelength * frequency = 2.2 * 450 = 990 m/s